Given an optimal risky portfolio with expected return of 12% and standard deviation of 26% and a risk free rate of 3%, what is the slope of the best feasible CAL?

Group of answer choices

0.64

0.14

0.08

0.35

0.36
Answer #1

slope of best feasible CAL= shaupe ratio = E(R) - Rf where Rf = 30%. 6- 26% ECR) = 12 % s slope = 12-3 26 = 0.35 Coption o)